OpenStudy (anonymous):

what is the 10th term of the sequence 81, 27, 9... a. 1/729 b. 1/243 c. 1/81 d. 1/810

OpenStudy (cwrw238):

this is a geometric sequence where each term is multiplied by a given number called a common ratio to get the next term can you see the pattern?

OpenStudy (cwrw238):

how do you get 27 from 81 and 9 from 27?

OpenStudy (cwrw238):

27 = x*81 what is x?

OpenStudy (cwrw238):

x = 28/81 = ?

OpenStudy (cwrw238):

* thats 27/81

OpenStudy (cwrw238):

you can simplify that by dividing 27 by 27 and 81 by 27;-